OpenStudy (anonymous):

Use the laws of exponents to simplify. 10^5/6 / 10^3/6

OpenStudy (anonymous):

add the exponents to get \[10^{\frac{8}{6}}\] \[10^{\frac{4}{3}}\]

OpenStudy (anonymous):

oh wait you are dividing!

OpenStudy (anonymous):

its a fraction

OpenStudy (anonymous):

in that case subtract to get \[10^{\frac{2}{6}}\] \[10^{\frac{1}{3}}\]

OpenStudy (anonymous):

your only real job here was to take \[\frac{5}{6}-\frac{3}{6}=\frac{2}{6}=\frac{1}{3}\]

jimthompson5910 (jim_thompson5910):

Subtract the exponents: \[\frac{5}{6}-\frac{3}{6}=\frac{5-3}{6}=\frac{2}{6}=\frac{1}{3}\] So the answer is \[10^{\frac{1}{3}}\]